OpenStudy (anonymous):

The defining equations for the function f is f(x) = 3x^2 + 1. The defining equation for the function g is g(x) = 2x-3. Then the defining equation for f \circle g (x) is:

OpenStudy (anonymous):

10x^2-5x-5 12x^2-36x+28 6x^2 + 28 5x^3 - 12x^2 + 9x - 1

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I dont understand what they mean

OpenStudy (anonymous):

work out f(g(x)) Begin by substituting the expression for g(x) into above f(g(x))

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

@phi

OpenStudy (phi):

f \circle g (x) means wherever you see an "x" in f(x), replace it with g(x) f(x) = 3x^2 + 1 replace x with g(x) f( g(x) ) = 3 (g(x))^2 + 1 they want you to "expand" the right side, by using the definition of g(x) they tell you g(x) = 2x-3 use that in the right side of f( g(x) ) = 3 (g(x))^2 + 1 f( g(x) ) = 3 (2x-3)^2 + 1 they may want you to expand that... it depends if it matches your answer choices.. and looking at them, it looks like you have to expand it. the first step if figure out (2x-3)^2 which is (2x-3)*(2x-3) can you do that ?